A heavy flywheel of inertia 0.5kgm^2 is put on a horizontal axis of a radius 2cm which has negligible mass. if a force of 50N is applied tangentially to the axle, what is the torque that rotates this flywheel if friction is neglected. ii. angular acceleration and ii. the angular velocity of the flywheel after 15s from rest
1
Expert's answer
2019-05-27T10:46:27-0400
"\tau=F\times l=50 \times 0.02=1 N\times m""\alpha=\tau/J=1/0.5=2 rad/sec^2""\omega=\omega_0+\alpha \times t=0+2\times15=30 rad/sec"